# Rule-Based Error Detection and Correction to Operationalize Movement Trajectory Classification

### v4

@@ 2 Background
[…]
@@ 3 Error Detection and Correction Rules
In this paper, we shall assume a set $\mathcal{O}$ of operational sequences for which there is ground truth available after model training. This set can be the set of training data, a subset, or a superset. Later, in our experiments, we look at cases where $\mathcal{O}=\mathcal{T}$ and $\mathcal{T}\subseteq\mathcal{O}$ however these are not requirements as our results are based on model performance on $\mathcal{O}$ and we envision use-cases where $\mathcal{O}$ is significantly different from $\mathcal{T}$. On these samples, for each class $i$, the model ($f_{\theta}$) returns class $i$ for $N_{i}$ of the samples, and for each class $i$ we have the number of true positives $TP_{i}$, false positives $FP_{i}$, true negatives $TN_{i}$, and false negatives $FN_{i}$. We have precision $P_{i}=TP_{i}/N_{i}=TP_{i}/(TP_{i}+FP_{i})$, recall $R_{i}=TP_{i}/(TP_{i}+FN_{i})$, and prior of predicting class $i$: $\mathbf{\mathcal{P}}_{i}=N_{i}/N$.
Language. We use a simple first-order language where samples are represented by constant symbols ($\omega$). We define set $C$ of $m$ “condition” unary predicates $cond_{1},\ldots,cond_{m}$ associated with each sample that can be either true or false - these are conditions that can be thought of as potentially leading to failure (but our learning algorithms will identify which ones lead to failure for a given prediction). We also define unary predicates for each class $i$: $pred_{i}$, $corr_{i}$, and $error$ defined below.
@@
  • $error$: This predicate is true if and only if an EDCR rule concludes there is an error in the model’s prediction.

### v5

@@ 2 Background
Movement Trajectory Classification Problem. The problem of classifying movement trajectories has been studied in the literature Zheng et al. (2008); Wang et al. (2017); Simoncini et al. (2018); Dabiri and Heaslip (2018); Kim et al. (2022) and we shall refer to it as the movement trajectory classification problem (MTCP). We also note that this line of work differs from and is complementary to trajectory generation Bavikadi et al. (2024); Janner et al. (2021); Chen et al. (2021); Itkina and Kochenderfer (2022) which does not seek to identify the mode of transportation. An MTCP instance is defined as given a sequence of GPS points, $\omega$, assign one of $n$ movement class from $\mathcal{C}$ which is often defined Kim et al. (2022); Dabiri and Heaslip (2018) as $\mathcal{C}=\{\textsf{walk},\textsf{bike},\textsf{bus},\textsf{drive},\textsf{train}\}$. The current paradigm for the MTCP problem is to create a neural model $f_{\theta}$ that maps sequences to movement classes using a set of weights, $\theta$. In this approach traditional methods (i.e., gradient descent) are used to find a set of parameters such that a loss function is minimized based on some training set $\mathcal{T}$ (where each sample $\omega\in\mathcal{T}$ is associated with a ground truth class $gt(\omega)$). Formally: $\arg\min_{\theta}\mathbb{E}_{\omega\in\mathcal{T}}\mathit{Loss}(f_{\theta}(\omega),gt(\omega))$. Within this paradigm, several approaches have been proposed. Most notably a CNN-based architecture Dabiri and Heaslip (2018) and the current state-of-the-art approach known as Long-term Recurrent Convolutional Network (LRCN) Kim et al. (2022) which combines lower CNN layers and upper LSTM layers - both of which we use as baselines in addition to an extension of LRCN that uses an additional attention head (LRCNa).
@@ 3 Error Detection and Correction Rules
In this paper, we shall assume a set $\mathcal{O}$ of operational sequences for which there is ground truth available after model training. This set can be the set of training data, a subset, or a superset. Later, in our experiments, we look at cases where $\mathcal{O}=\mathcal{T}$ and $\mathcal{T}\subseteq\mathcal{O}$ - however these are not requirements as our results are based on model performance on $\mathcal{O}$ - and we envision use-cases where $\mathcal{O}$ is significantly different from $\mathcal{T}$. On these samples, for each class $i$, the model ($f_{\theta}$) returns class $i$ for $N_{i}$ of the samples, and for each class $i$ we have the number of true positives $TP_{i}$, false positives $FP_{i}$, true negatives $TN_{i}$, and false negatives $FN_{i}$. We have precision $P_{i}=TP_{i}/N_{i}=TP_{i}/(TP_{i}+FP_{i})$, recall $R_{i}=TP_{i}/(TP_{i}+FN_{i})$, and prior of predicting class $i$: $\mathbf{\mathcal{P}}_{i}=N_{i}/N$.
Language. We use a simple first-order language where samples are represented by constant symbols ($\omega$). We define set $C$ of $m$ “condition” unary predicates $cond_{1},\ldots,cond_{m}$ associated with each sample that can be either true or false - these are conditions that can be thought of as potentially leading to failure (but our learning algorithms will identify which ones lead to failure for a given prediction). We also define unary predicates for each class $i$: $pred_{i}$, $corr_{i}$, and errorerror defined below.
@@
  • errorerror: This predicate is true if and only if an EDCR rule concludes there is an error in the model’s prediction.